# How to Fine-Tune a Reasoning Model? A Teacher–Student Cooperation Framework to Synthesize Student-Consistent SFT Data

### 23 Mar 2026 v1

@@ 1 Introduction
Recently, major AI companies have increasingly introduced reasoning models as flagship large language models (LLMs) [openai_o1, dsr1, qwen3]. These models generate responses that bifurcate into a thinking content for fine-grained reasoning and a final answer content that delivers the solution. By introducing explicit thinking processes, LLMs have achieved new breakthroughs in reasoning-intensive tasks such as code generation. [lcb_pro, ojbench].
@@ 3.2 Datasets
Our primary experiments focused on the code generation task, with mathematics and science question answering tasks used as auxiliary tests for out-of-domain evaluation. For code generation, we evaluated LiveCodeBench-V5 (2024/08/01–2025/02/01) and LiveCodeBench-V6 (2025/02/01–2025/05/01) [lcb], as well as LiveCodeBench-Pro [lcb_pro] and OJBench [ojbench]. For the auxiliary tasks, we evaluated AIME-2024, AIME-2025, and OlympiadBench [olympiadbench] for mathematics, along with GPQA [gpqa] for science. For datasets containing multimodal data, only text-only parts were used.
@@
LCB-V5 LCB-V6 LCB-Pro OJBench GPQA AIME2024 AIME2025 OlympiadBench
@@
+ Teacher-Score 54.55 48.00 26.06 18.32 60.61 75.42 66.77 69.53
+ Teacher-Reference 51.50 43.43 19.26 14.22 60.61 71.98 62.29 67.80
@@ 3.4 Compared Methods
We compared three categories of SFT data construction settings: student-driven, teacher–student collaborative, and teacher-driven data synthesis. The student-driven setting included Teacher-Reference (i.e., Self-Distillation), where the teacher first generated a reference answer and the student generated the thinking and final answer based on the reference answer [SDFT], and Teacher-Score (i.e., Reject-Sampling), where the student generated multiple candidates and the teacher scored and selected the best one. In our experiments, five candidates were generated for selection. The teacher–student collaborative setting included Teacher-Answer, in which the student generated thinking content followed by the teacher producing the final answer, and Teacher-Think, which assigned thinking generation to the teacher and answer generation to the student. The teacher-driven setting included Teacher-Only [seq_distill], where all training samples were fully generated by the teacher, as well as Teacher-Mix, which mixed samples generated by the teacher and the student in a 1:1 ratio [lmitKD].
@@ 3.5 Experimental Results
Table 1 presents a comparison between TESSY and other SFT data synthesis approaches. Although training in synthetic data constructed by all other methods led to varying degrees of degradation in code generation performance, TESSY consistently improved Qwen-8B in all four datasets. Specifically, TESSY increased the performance of Qwen-8B by 7.78%, 5.85%, 11.34%, and 6.68% on the three LiveCodeBench datasets and OJBench, respectively. In contrast, only the Teacher-Score method provided a marginal improvement of 0.71% on LiveCodeBench-Pro. Meanwhile, the widely used Teacher-Only approach, even after fully training up to 80K samples for 9 epochs, still suffered a performance drop of up to 10.02% in OJBench.
Observations on baselines indicate that, in general, the greater the teacher’s participation in data synthesis, the greater the drop in student performance. For example, on OJBench, the fully teacher-generated method Teacher-Only caused a decrease of 10.02%. However, although reducing teacher involvement in data generation helps maintain model stability, it does not necessarily lead to performance gains, as exemplified by Teacher-Score. In contrast, TESSY, despite using up to 77.65% teacher-generated tokens as shown in Figure 17, did not cause any performance degradation and instead consistently improved the student model on in-domain datasets.
Furthermore, TESSY demonstrated potential for generalization to out-of-domain datasets. On AIME2024 and AIME2025, training with the data synthesized by TESSY improved Qwen3-8B by 3.75% and 0.93%, respectively, while preserving comparable performance on GPQA and OlympiadBench. In contrast, other methods relying on teacher-generated data led to varying degrees of performance degradation. For example, Teacher-Only caused performance drops of 4.35%, 2.71%, and 2.05% on GPQA, AIME2025, and OlympiadBench, respectively.
@@ 5 Related Work
Training student models using data synthesized by a teacher is a widely adopted approach [seq_distill, lmitKD]. However, due to differences in capacity and style, a stronger teacher does not necessarily produce a stronger student [small_struggle]. Some methods adapt the teacher to a student-friendly distribution, but retraining teachers with hundreds of billions of parameters is often prohibitively expensive [student_friendly_teacher, a_good_learner]. Self-distillation mitigates the distribution mismatch by having the student rewrite teacher-generated references [SDFT, SDFT_plus], but this process can introduce shortcut behaviors in reasoning models. The closest works are on-policy distillation [on_policy_distill], where the teacher supervises student-generated data, and methods like AdaSwitch further improve this via alternating generation [AdaSwitch]. However, training models with different vocabularies remains a challenge. Moreover, unlike these online approaches, our work focuses on synthesizing offline SFT data that efficiently transfers knowledge across models and tasks, enabling practical and open use.
@@ 6 Conclusion
To address the challenge of further improving reasoning models through supervised fine-tuning, we propose a teacher–student cooperation data synthesis framework, in which teacher and student models alternately synthesize responses for training. By involving the student model in the data synthesis, our approach preserves essential reasoning information from the teacher while aligning the synthesized data distribution with the inherent style of the student. In the future, we will further refine the framework to enable a more accurate and efficient identification of style and capability boundaries. In addition, we plan to extend the proposed cooperative paradigm to a wider range of tasks and heterogeneous model settings, exploring its potential for scalable multi-model collaboration.
@@
Appendix A Appendix
@@
A.1 Model Selection for Final Answer Generation
@@
A.2 Comparison of System Prompt Isolation
@@ A.3 Performance Over Epochs
A.4 Comparison of Average Token Counts
@@ A.5 Performance of LoRA
A.6 Changes in Model Output Similarity
@@
A.7 Training Loss Curve

### 21 Apr 2026 v2

@@ 1 Introduction
Recently, major AI companies have increasingly introduced reasoning models as flagship large language models (LLMs) [openai_o1, dsr1, qwen3]. These models generate responses that bifurcate into a thinking content for fine-grained reasoning and a final answer content that delivers the solution. By introducing explicit thinking processes, LLMs have achieved new breakthroughs in reasoning-intensive tasks such as code generation [sun2024survey, lcb_pro, ojbench].
@@ 3.2 Datasets
Our primary experiments focused on the code generation task, with mathematics and science question answering tasks used as auxiliary tests for out-of-domain evaluation. For code generation, we evaluated LiveCodeBench-V5 (2024/08/01–2025/02/01) and LiveCodeBench-V6 (2025/02/01–2025/05/01) [lcb], as well as LiveCodeBench-Pro [lcb_pro] and OJBench [ojbench]. For the auxiliary tasks, we evaluated AIME-2024, AIME-2025, and OlympiadBench [olympiadbench] for mathematics, along with GPQA-Diamond [gpqa] for science. For datasets containing multimodal data, only text-only parts were used.
@@
LCB-V5 LCB-V6 LCB-Pro OJBench GPQA-Diamond AIME2024 AIME2025 OlympiadBench
@@
+ Reject-Sampling 54.55 48.00 26.06 18.32 60.61 75.42 66.77 69.53
+ Self-Distillation 51.50 43.43 19.26 14.22 60.61 71.98 62.29 67.80
@@ 3.4 Compared Methods
We compared three categories of SFT data construction settings: student-driven, teacher–student collaborative, and teacher-driven data synthesis. The student-driven setting included Self-Distillation, where the teacher first generated a reference answer and the student generated the thinking and final answer based on the reference answer [SDFT], and Reject-Sampling, where the student generated multiple candidates and the teacher scored and selected the best one. In our experiments, five candidates were generated for selection. The teacher–student collaborative setting included Teacher-Answer, in which the student generated thinking content followed by the teacher producing the final answer, and Teacher-Think, which assigned thinking generation to the teacher and answer generation to the student. The teacher-driven setting included Teacher-Only [seq_distill], where all training samples were fully generated by the teacher, as well as Teacher-Mix, which mixed samples generated by the teacher and the student in a 1:1 ratio [lmitKD].
@@ 3.5 Experimental Results
Table 1 presents a comparison between TESSY and other SFT data synthesis approaches. Although training in synthetic data constructed by all other methods led to varying degrees of degradation in code generation performance, TESSY consistently improved Qwen-8B in all four datasets. Specifically, TESSY increased the performance of Qwen-8B by 7.78%, 5.85%, 11.34%, and 6.68% on the three LiveCodeBench datasets and OJBench, respectively. In contrast, only the Reject-Sampling method provided a marginal improvement of 0.71% on LiveCodeBench-Pro. Meanwhile, the widely used Teacher-Only approach, even after fully training up to 80K samples for 9 epochs, still suffered a performance drop of up to 10.02% in OJBench.
Observations on baselines indicate that, in general, the greater the teacher’s participation in data synthesis, the greater the drop in student performance. For example, on OJBench, the fully teacher-generated method Teacher-Only caused a decrease of 10.02%. However, although reducing teacher involvement in data generation helps maintain model stability, it does not necessarily lead to performance gains, as exemplified by Reject-Sampling. Notably, although Self-Distillation is widely regarded as a promising approach to mitigate distribution mismatch in training data, our experiments reveal that it still incurs substantial performance degradation, consistent with the findings reported in [sd_degrade_reason]. In contrast, TESSY, despite using up to 77.65% teacher-generated tokens as shown in Figure 20, did not cause any performance degradation and instead consistently improved the student model on in-domain datasets.
Furthermore, TESSY demonstrated potential for generalization to out-of-domain datasets. On AIME2024 and AIME2025, training with the data synthesized by TESSY improved Qwen3-8B by 3.75% and 0.93%, respectively, while preserving comparable performance on GPQA-Diamond and OlympiadBench. In contrast, other methods relying on teacher-generated data led to varying degrees of performance degradation. For example, Teacher-Only caused performance drops of 4.35%, 2.71%, and 2.05% on GPQA-Diamond, AIME2025, and OlympiadBench, respectively.
@@ 5 Related Work
Training student models using data synthesized by a teacher is a widely adopted approach [seq_distill, lmitKD]. However, due to differences in capacity and style, a stronger teacher does not necessarily produce a stronger student [small_struggle]. Some methods adapt the teacher to a student-friendly distribution, but retraining teachers with hundreds of billions of parameters is often prohibitively expensive [student_friendly_teacher, a_good_learner]. Self-distillation mitigates the distribution mismatch by having the student rewrite teacher-generated references [SDFT, SDFT_plus], but this process can introduce shortcut behaviors in reasoning models. The closest works are on-policy distillation [on_policy_distill], where the teacher supervises student-generated data, and methods Speculative Distillation and AdaSwitch further improve this via alternating generation [speculative_distill, AdaSwitch]. However, training models with different vocabularies remains a challenge. Moreover, unlike these online approaches, our work focuses on synthesizing offline SFT data that efficiently transfers knowledge across models and tasks, enabling practical and open use.
@@ 6 Conclusion
To address the challenge of further improving reasoning models through supervised fine-tuning, we propose a teacher–student cooperation data synthesis framework, in which teacher and student models alternately synthesize responses for training. By involving the student model in the data synthesis, our approach preserves essential reasoning information from the teacher while aligning the synthesized data distribution with the inherent style of the student. In the future, we will further refine the framework to enable a more accurate and efficient identification of style and capability boundaries. Due to engineering constraints, TESSY’s synthesis speed remains below its theoretical limit. We will optimize the system implementation and explore acceleration techniques [medusa, branch_decoder, pipeline_decoder]. In addition, we plan to extend the proposed cooperative paradigm to a wider range of tasks and heterogeneous model settings, exploring its potential for scalable multi-model collaboration.
@@
Appendix A More Comparisons
Table 2: Performance comparison using the DeepSeek-R1-Distill-Llama-8B as the student model.
LCB-V5 LCB-V6 LCB-Pro OJBench
DeepSeek-R1-Distill-Llama-8B 34.73 33.71 17.56 5.82
+ Self-Distillation 19.76 24.00 8.07 5.82
+ Teacher-Only 38.32 36.57 24.22 11.42
+ TESSY 52.10 41.71 26.20 12.50
A.1 Evaluation Beyond the Qwen3 Family
In Table 2, we evaluated the performance of TESSY when using DeepSeek-R1-Distill-Llama-8B  [^1] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B as the student model (with GPT-OSS-120B as the teacher and trained on 40K samples). The results show that TESSY consistently improves the reasoning capability of the student model, whereas self-distillation still leads to severe catastrophic forgetting.
In particular, although the Teacher-Only method performed significantly worse than TESSY, unlike the results observed in Qwen3-8B in Table 1, it improved code reasoning performance in DeepSeek-R1-Distill-Llama-8B on several benchmarks. This discrepancy may be attributed to the fact that Qwen3-8B is a more thoroughly post-trained reasoning model. As a result, introducing distributionally mismatched SFT data to a relatively well-trained model leads to more pronounced catastrophic forgetting.
Figure 9: In-domain evaluation on the math reasoning task (AIME2025), where GPT-OSS-120B and DeepSeek-R1-Distill-Llama-8B were used as the teacher and student models, respectively.
Figure 10: In-domain evaluation on the science reasoning task (GPQA-Diamond), where GPT-OSS-120B and Qwen3-8B were used as the teacher and student models, respectively.
A.2 Evaluation on Math Reasoning Tasks
In Table 1, we reported the out-of-domain performance on mathematical reasoning tasks when using pure code data as the training set. To further validate the effectiveness of TESSY on other reasoning tasks, we constructed an SFT dataset using open-source math questions in this section for evaluation. However, we found that Qwen3 family models are already sufficiently trained on these datasets, such that Qwen3-8B even outperforms the teacher model GPT-OSS-120B on the training set. To provide a more faithful evaluation, we therefore used DeepSeek-R1-Distill-Llama-8B instead of Qwen3-8B for comparison.
We randomly sampled 12K questions from DeepMath [deepmath] for data synthesis, and reported the results in Figure 10. For both TESSY and the baselines, we applied rejection sampling to filter the data and retained only samples with correct answers. Consistent with previous findings, on AIME2025, TESSY outperforms Teacher-Only by up to 4.28%, while the Self-Distillation method leads to severe degradation in reasoning performance. We have released the training data synthesized by TESSY  [^2] https://huggingface.co/datasets/CoopReason/TESSY-Math-12K.
A.3 Evaluation on Science Reasoning Task
We further conducted in-domain evaluation of TESSY on GPQA-Diamond. We randomly selected 3K questions from SuperGPQA [supergpqa] in Physics, Chemistry, and Biology to construct the SFT dataset. For both TESSY and the baselines, we applied rejection sampling to filter the data and retained only samples with correct answers. As shown in Figure 10, TESSY improves the performance of Qwen3-8B by 2.72%, whereas Teacher-Only leads to a 5.11% performance degradation. We have released the training data synthesized by TESSY  [^3] https://huggingface.co/datasets/CoopReason/TESSY-SuperGPQA-3K.
Figure 11: Comparison of Teacher-Restyle baseline, where we prompt teacher model to generate response as the student style.
A.4 Comparison of Teacher Output Style Control
In Figure 12, we evaluated a baseline termed Teacher-Restyle, which conditions the teacher model with a prompt that specifies the desired style of the final answer of Qwen3-8B. Specifically, we provided an exemplar response generated by Qwen3-8B to instruct the teacher to produce reasoning traces aligned with this target output style. However, although human observation suggests that the final answers generated by the teacher model exhibit a style highly similar to that of the student model, Teacher-Restyle still leads to severe catastrophic forgetting in Qwen3-8B. This may be because the long-form thinking part is difficult to control in style through simple prompting.
A.5 Performance of LoRA
@@
A.6 Comparison of System Prompt Isolation
A.7 Model Selection for Final Answer Generation
@@
Table 3: Comparison of training performance using TESSY synthetic data with boundary predictors trained on different dataset sizes. Different from Table 1 (80K samples), we used 40K synthesized samples here to reduce computational cost for comparison.
LCB-V5 LCB-V6 LCB-Pro OJBench
Qwen3-8B 55.09 49.58 25.35 18.75
+ TESSY (Predictors trained on 500 samples) 59.58 54.00 35.13 23.98
+ TESSY (Predictors trained on 100K samples) 60.68 55.15 35.84 26.51
A.8 Impact of Predictor Training Size
In the main experiments, we used the teacher model to automatically construct 100K training samples for both the teacher and student boundary predictors. It is worth noting that this scale is over-provisioned to better distill the teacher model’s capability. In practice, boundary prediction is not a complex task, and a relatively small amount of data is sufficient for Qwen3-0.6B-Base to achieve strong performance.
To verify the impact of training data size for the predictor on the TESSY framework, in Table 3, we compared the main setting using 100K training samples with a smaller setting using only 500 samples for training the predictors. It can be observed that, while 100K training samples allow boundary predictors to better support TESSY in synthesizing effective training data, even with only 500 samples, TESSY still functions properly, avoids catastrophic forgetting, and improves the performance of Qwen3-8B. In addition, the training data for boundary predictors is constructed fully automatically, requiring no additional human effort, and the computational cost of scaling the dataset is negligible. Therefore, in practice, we recommend using the over-provisioned training data as adopted in the main experiments of this work.
Another finding is that TESSY is already effective with only 40K training samples. In the main experiments, we used up to 80K samples and trained for 9 epochs to avoid the possibility that baselines could mitigate distribution mismatch issues under increased computational scaling. However, as shown in Table 3, we find that even when baselines are trained with twice the amount of data used by TESSY, they still fail to outperform TESSY. The effect of the number of training epochs can be found in Section B.1.
Appendix B More Analysis
B.1 Performance Over Epochs
@@
B.2 Comparison of Average Token Counts
@@
B.3 Changes in Model Output Similarity
@@
B.4 Training Loss Curve